### v1

# Identification and Control of Neutral Anyons

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: Email: jia__\__li@brown.edu
Identification and Control of Neutral Anyons
Ron Q. Nguyen1    Naiyuan J. Zhang1    Navketan Batra1,2    Xiaoxue Liu3    Kenji Watanabe4    Takashi Taniguchi5    D. E. Feldman1,2    J.I.A. Li1 1Department of Physics, Brown University, Providence, Rhode Island 02912, USA 2Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA 3Tsung-Dao Lee Institute and School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China 4Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 5Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(October 31, 2024)
Beyond the well-known fermions and bosons, anyons—an exotic class of particles—emerge in the fractional quantum Hall effect and exhibit fractional quantum statistics. Anyons can be categorized by their charge, with extensive research focused on those carrying fractional charge, while charge-neutral anyons in 2D electron liquids remain largely unexplored. Here, we introduce bilayer excitons as a new pathway to realizing charge-neutral anyons. By pairing quasiparticles and quasiholes from Laughlin states, we report bilayer excitons that obey fractional quantum statistics. Through layer-asymmetric field-effect doping, we achieve precise control of the anyon population, stabilizing anyonic dipoles at temperatures below the charge gap. Furthermore, we investigate neutral anyons in even-denominator FQHE states, which are likely described by non-Abelian wavefunctions. These findings open the door to exploring non-Abelian statistics in neutral anyons, with the potential to reshape future research in topological quantum phases.
Figure 1: Exciton pairing between a pair of Laughlin states. (a) Schematic of the double-layer graphene device used in this geometry to stabilize spatially indirect exciton pairing between quasiparticles and quasiholes across two graphene layers. (b) Schematic of the edgeless Corbino geometry used for transport measurements in parallel flow (PF), counterflow, and drag configurations. (c) Schematic diagram showing FQHE within the CF model, where two flux attachments are included in the CF construction. Blue and red solid lines denote integer values for $\nu^{\ast}_{1}$ and $\nu^{\ast}_{2}$, respectively  Li2019pairing ; Liu2019interlayer . At the intersection of these lines, both layers simultaneously develop a robust charge gap, leading to an incompressible state across the quantum Hall bilayer. (d-e) Transport response along $\nu_{1}^{\ast}=1$, at (d) $B=12$ T and (e) $B=28$ T. The top, middle, and bottom panels show PF conductance $G_{\text{PF}}$, counterflow conductance $G_{\text{CF}}$, and drag ratio $I_{\text{drag}}/I_{\text{drive}}$, respectively. Inset of each panel shows schematic measurement configurations in the parallel flow, counterflow, and drag geometries. Blue, green, and red circles mark different ground state orders.
Figure 2: Pairing between quasiparticle excitations. (a) Arrhenius plot of bulk conductance measured in the parallel flow $G_{\text{PF}}$ (black) and counterflow $G_{\text{CF}}$ (blue) configurations at $\nu_{1}=\nu_{2}=1/3$ and $B$ = 28 T. Inset shows thermally excitated quasiparticles and quasiholes across the charge gap of the Laughlin state. (b) $B$-dependence of counterflow conductance $G_{\text{CF}}$ in logarithmic scale measured at $\nu_{1}=\nu_{2}=1/3$ (Black circles), $2/5$ (blue circles), and $3/7$ (Red circles).
Figure 3: Doping neutral anyons into the Laughlin state. (a) Controlling the population of neutral anyons through layer-asymmetric doping, which is achieved by varying Landau level fillings of layer 1 and 2 by ±$\pm\delta$. This generates equal number of quasiparticles and quasiholes across two graphene layers. (b) Parallel flow conductance $G_{\text{PF}}$ (left), drag ratio $I_{\text{drag}}/I_{\text{drive}}$ (middle), and schematic diagram (right) summarizing the transport response across the $\nu_{\text{total}}$-$\Delta\nu$ map around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. Near $\Delta\nu$=0, $I_{\text{drag}}$ and $I_{\text{drive}}$ both vanish, resulting in ill-defined drag ratio masked by gray-shaded areas. Black arrows in the right panel mark the extend of the robust charge gap along the direction of $\Delta\nu$ and $\nu_{\text{total}}$. (c) $G_{\text{PF}}$, $G_{\text{CF}}$ (top) and $I_{\text{drag}}/I_{\text{drive}}$ (bottom) as a function of $\Delta\nu$ measured at $\nu_{\text{total}}=2/3$. (d) Arrhenius plot of $G_{\text{PF}}$ (black) and $G_{\text{CF}}$ (blue) measured at $\Delta\nu=0.06$. (e) $G_{\text{PF}}$, $G_{\text{CF}}$ as a function of varying $\nu_{1}$ while $\nu_{2}$ is fixed at $1/3$.
Figure 4: Non-Abelian neutral mode at $\nu_{1}=\nu_{2}=3/8$. (a) A detailed view of the $\nu_{\text{total}}$-$\Delta\nu$ phase space, where black and gray lines indicate CF states with two and four flux attachments, respectively. (b) Parallel flow conductance $G_{\text{PF}}$, (c) counterflow conductance $G_{\text{CF}}$, and (d) drag ratio $I_{\text{drag}}/I_{\text{drive}}$ as a function of varying $\nu_{\text{total}}$ measured at $B=31$ T. Vertical arrows highlight different types of charge-neutral anyons emerging at locations marked by red, blue, and green circles in panel (a). (e) $G_{\text{PF}}$ as a function of $\nu_{\text{total}}$ and $B$, revealing a series of $B$-induced transitions in the FQHE states, marked by horizontal black solid lines. (f) $G_{\text{PF}}$ measured across a portion of the phase space encircled by the red dashed line in panel (a). Vanishing $G_{\text{PF}}$ defines incompressible states marked by black and gray circles in panel (a).
Quantum statistics underpins our understanding of various quantum phases of matter, a cornerstone of condensed matter research. In an electronic system, fermionic statistics are key to describing various quantum phases emerging from a low-temperature electronic systems  Luttinger1960fermi . In contrast, excitons, which are bound states of electrons and holes, provide a pathway to explore bosonic ground states in solid state systems, such as Bose-Einstein condensation  Blatt1962exciton ; Lozovik1975exciton ; Lozovik1976exciton . An even more exotic class of quantum particles, known as anyons, emerges in the context of the fractional quantum Hall effect  Tsui1982FQHE ; Wilczek1982anyon ; Wilczek1982fractional ; Arovas1984fractional ; review-FH ; Picciotto1997 ; Reznikov1999 ; Dolev2008 . Driven by the rich interplay between electron correlation and topology within a partially occupied Landau level (LL), quasiparticle excitations are shown to carry a fractional multiple of the electron’s charge. As a defining characteristic of the FQHE, this fractional charge plays a pivotal role in shaping its transport properties, including the hallmark signature of quantized Hall conductance at fractional values of $e^{2}/h$  Tsui1982FQHE .
Although fractional statistics are often associated with charge-carrying particles  Wilczek1982anyon ; Wilczek1982fractional ; Arovas1984fractional ; Picciotto1997 ; Reznikov1999 ; Dolev2008 , neutral particles can also exhibit anyonic behavior. A unique class of neutral anyons has been discussed in the context of non-Abelian wavefunctions. A notable example is the chiral Majorana mode, which obeys non-Abelian Ising statistics and enables the braiding operations in distinct fusion sectors, a critical feature for topological qubits  Nayak.08 . Recent experiments have leveraged the thermal conductance of such neutral anyons as a key fingerprint to distinguish between candidate topological orders, such as the particle-hole Pfaffian (PH-Pf) and anti-Pfaffian (A-Pf) states  Banerjee2017a ; Banerjee2017b . Unlike quantum phases governed by Coulomb interactions between charged particles, excitonic systems follow fundamentally different dynamics due to their charge neutrality, opening new avenues for exploring exotic quantum states. The charge neutrality also decouples neutral anyons from environmental interference, enhancing their robustness. The experimental realization of neutral anyons, therefore, holds great potential to broaden our understanding of quantum phases of matter with exotic topology, presenting a new frontier for condensed matter physics.
Charge-neutral anyons, by definition, do not carry net charge, rendering conventional charge transport techniques ineffective in probing their behavior. This limitation has driven the development of new methods to study these exotic particles. We report a novel approach for stabilizing, identifying, and controlling charge-neutral anyons in quantum Hall bilayer systems, achieved through excitonic pairing between bulk quasiparticles and quasiholes, both of which exhibit anyonic statistics. This unlocks the intriguing possibility of an anyonic exciton. Using field-effect doping, we precisely control the population of these neutral anyons by introducing equal numbers of bulk quasiparticles and quasiholes into two graphene layers. This induces a transition from a state, devoid of bilayer excitons, to a distinct phase where stable excitonic anyons persist even at temperatures significantly below the charge gap. Our findings highlight a novel ground state consisting of interlayer “anyon dipoles” coexisting with the FQHE in both layers.
Amongst a multitude of neutral anyons observed, we particularly highlight those associated with an even-denominator state, where the likely ground state order is non-Abelian in nature. This leads to neutral excitations obeying non-Abelian statistics. Our findings open new avenues for exploring charge-neutral anyon with exotic properties, providing new insights into the intricate interplay between correlation, topology, and charge fractionalization.
The quantum Hall bilayer system confines charge carriers in closely separated but electrically isolated graphene layers, as illustrated in Fig. 1a  Li2017superfluid ; Liu2017superfluid ; Li2019pairing ; Liu2022crossover . To probe bulk quasiparticles and quasiholes, we pattern the bilayer system into the edgeless Corbino geometry  Nandi2012exciton ; Eisenstein2014 ; Li2019pairing ; Zeng2023solid ; Zhang2024fractionalexciton , which enables three distinct electrical transport configurations (Fig. 1b, also see Methods Section for more detailed discussion). The bulk conductance measurement in the parallel flow (PF) geometry probes the charge gap of the bilayer system, while the counterflow and drag geometries serve as direct probes of charge neutral-modes with non-zero dipole moment, such as interlayer excitons.
When subjected to an out-of-plane magnetic field, the LL filling of each graphene layer, $\nu_{1}$ and $\nu_{2}$, can be independently controlled by adjusting the voltage bias on the top and bottom gate electrodes, $V_{1}$ and $V_{2}$. This configuration allows access to a low-temperature phase space defined by the total filling $\nu_{\text{total}}=\nu_{1}+\nu_{2}$ and the layer imbalance $\Delta\nu=\nu_{1}-\nu_{2}$. We focus on a region of the phase space dominated by the conventional Jain sequence of FQHE states, whose locations are indicated by black circles in Fig. 1c. At these locations, layer 1 is occupied by a Laughlin state at $\nu_{1}=1/3$, while layer 2 is tuned across a sequence of Jain states characterized by $\nu_{2}^{\ast}=\nu_{2}/(1-2\nu_{2})=N$, where $N\in\mathbb{Z}$ defines the sequence. In this work, we use $\nu_{i}^{\ast}$ to denote composite fermion (CF) effective filling $\nu^{\ast}$ in layer $i=1,2$  Jain.03 ; Eisenstein1990FQHE . These CFs are constructed with intralayer flux attachments only, distinguishing them from previous reports of two-component CFs  Li2019pairing ; Zhang2024fractionalexciton .
Fig. 1d plots transport measurement across this sequence of Jain states at $B=12$ T. As shown in the top panel, the PF geometry reveals vanishing bulk conductance, confirming the presence of robust charge gaps at the expected positions of Jain states. Interestingly, the insulating states in the PF geometry exhibits non-zero conductance in the counterflow measurement (middle panel of Fig. 1d), which also coincide with the perfect drag response (bottom panel of Fig. 1d). As a current bias is applied to the drive layer (layer 1), a counter-propagating current of equal magnitude is detected in the drag layer (layer 2), giving rise to a unity drag ratio $I_{\text{drag}}/I_{\text{drive}}=1$. These observations suggest the presence of neutral modes, likely arising from interlayer exciton pairing.
It is important to note that the Jain sequence of FQHE states, characterized by its hierarchical behavior shown in Fig. 1d, is described by well-defined wavefunctions localized within each layer. The absence of interlayer correlation in the CF construction raises an intriguing conundrum regarding the origin of the neutral mode. An interesting $B$-dependence offers an important clue. As shown in Fig. 1e, the neutral mode between a pair of Laughlin state at $\nu_{1}=\nu_{2}=1/3$ vanishes at $B=28$ T. At this large magnetic field, vanishing conductance is observed in both the PF and counterflow geometries (top and middle panels of Fig. 1e). In the drag geometry, both the drive and drag currents approach zero, leading to an ill-defined drag ratio, which is masked with the gray-shaded region in the bottom panel of Fig. 1e.
We propose that the neutral mode arises from exciton pairing between negatively charged quasiparticle in one layer and a positively charged quasihole in the other. As shown in the inset of Fig. 2a, quasiparticles and quasiholes are thermally generated across the charge gap. As such, their population is governed by the temperature in relation to the charge gap. In systems with a large charge gap, the neutral modes are suppressed at low temperatures, but are expected to re-emerge as the temperature increases. This hypothesis is validated by the temperature dependence of $G_{\text{PF}}$ and $G_{\text{CF}}$, measured at $B=28$ T and $\nu_{1}=\nu_{2}=1/3$, as shown in Fig. 2a. The activated behavior of $G_{\text{PF}}$ is directly linked to the charge gap, $\Delta_{\text{charge}}$, which represents the energy required to create thermally excited quasiparticles and quasiholes. Interestingly, the activated behavior of $G_{\text{CF}}$ reveals the energy required to generate a charge-neutral excitation, $\Delta_{\text{exciton}}$, which closely aligns with $\Delta_{\text{charge}}$. This provides direct evidence that both quasiparticles and the neutral mode are generated through the same thermal process.
It should be noted that the Jain states follow a hallmark hierarchical behavior. As the effective magnetic field decreases at larger effective filling, the charge gap is reduced, making it more preferable to generate quasiparticles and quasiholes. This has a direct impact on the transport response in the counterflow channel. As shown in the bottom panel of Fig. 1e, while the Laughlin state exhibits zero counterflow conductance at $B=28$ T, a perfect drag response is recovered as layer 2 is tuned to Jain states at larger effective fillings.
Along the same vein, Fig. 2b plots counterflow conductance $G_{\text{CF}}$ as a function of $B$ for three pairs of Jain states. The presence of charge-neutral modes, indicated by a non-zero $G_{\text{CF}}$, disappears at $\nu_{1}$ = $\nu_{2}$ = 1/3 after $B=15$ T, while persisting up to $B$ = 30 T at $\nu_{1}$ = $\nu_{2}$ = 2/5 and even higher $B$ at $\nu_{1}$ = $\nu_{2}=3/7$. This comparison establishes a general behavior: at each filling, the neutral mode is suppressed as the charge gap is enhanced with increasing $B$, while at each $B$, the neutral mode is enhanced as the charge gap is suppressed at larger effective filling. The direct correlation between the charge gap and neutral mode stability strongly supports the proposed pairing mechanism between quasiparticles.
Together, our observations provide strong evidence for neutral anyons composed of fractionally charged components. The anyonic nature stems from the fact that both the quasiparticles and quasiholes in the Laughlin state share the same braiding phase of $32\pi/$ 3. Consequently, the combined neutral exciton exhibits a braiding phase of $34\pi/$ 3. We note that the quantum statistics of the neutral mode can be tuned by selecting its parent Jain state. For instance, the neutral mode derived from $\nu_{1}=1/3$ and $\nu_{2}=2/3$ is expected to be bosonic, as the quasiparticles and quasiholes exhibit opposite braiding phase of ±$\pm 2\pi/3$ (see Fig. M1 and Method Section II for further details). This results in a robust ground state described by the Bose-Einstein condensation, which is evidenced by the perfect drag response observed at low temperature and $B=28$ T (Fig. 1e).
We note that the $B$-dependence of the neutral mode cannot be accounted for by the effective interlayer separation, $d/\ell_{B}$, a parameter typically used to describe the stability of the exciton condensate at a total filling of one  Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Liu2022crossover . Here, the magnetic length $\ell_{B}$ governs intralayer correlations, while the interlayer separation $d$ defines interlayer correlations. It is important to distinguish between the neutral anyons in Laughlin states, where the charge gap emerges solely from intralayer correlations, and the interlayer excitons of the (111) state Halperin1983 ; Wen1992 , a two-component ground state where interlayer coupling is essential. As a result, $d/\ell_{B}$ is inadequate for describing the properties of neutral anyons.
When quasiparticles are thermally generated, their corresponding neutral mode also behaves as a thermal excitation that disappears at low temperatures, as shown in Fig. 1d and Fig. 2a. In comparison, the presence of neutral anyons as a ground state feature would be highly unusual, potentially leading to exotic phases such as the anyonic superconductor  chen1989anyon ; lee1989anyon . In this work, we show that the low-temperature population of neutral anyons can be controlled through field-effect doping.
As illustrated in Fig. 3a, an equal population of quasiparticles and quasiholes can be generated across the two graphene layers by doping the bilayer to $\nu_{1}=1/3+\delta$ and $\nu_{2}=1/3-\delta$. As quasiparticles and quasiholes pair to form charge-neutral anyons, the charge gap remains robust against this form of doping, which describes a vertical line in phase space as a function of varying layer imbalance $\Delta\nu$, while keeping the total filling fraction $\nu_{\text{total}}$ constant.
The presence of doping-induced excitons is demonstrated by the transport properties around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$. Vanishing parallel flow conductance with $G_{\text{PF}}=0$ highlights a region with a fully developed charge gap, represented by dark blue in the left panel of Fig. 3b. While this region spans a range of $0.20$ in the direction of $\Delta\nu$, it covers a much narrower range of only $0.02$ in the direction of $\nu_{\text{total}}$. This directional dependence aligns well with the unique doping requirement for exciton formation. Tuning along other directions, such as varying $\nu_{\text{total}}$ while holding $\Delta\nu$ fixed, leads to mismatched quasiparticle and quasihole populations, resulting in unpaired charges and a lack of drag response.
Within the boundary of the charge gap, drag measurements reveal an intriguing crossover. In the region near $\Delta\nu$ $=0$, the bilayer sample exhibits vanishing conductance in both parallel flow and counterflow measurements, as illustrated in Fig. 3c. As both drive and drag currents are zero, the drag ratio is ill-defined, which is indicated by the gray-shaded area in Fig. 3b and c. As $\Delta\nu$ is detuned from zero, the counterflow conductance rises from zero while the parallel flow conductance remains vanishingly small. This leads to a regime characterized by a perfect drag response, represented by dark red in the middle and right panels of panel of Fig. 3b and indicated in the red-shaded area of Fig. 3c. These transport responses point to excitonic pairing, driven by the doping-induced population of quasiparticles and quasiholes.
In the presence of doping-induced excitons, the bulk conductance in the counterflow geometries no longer exhibits activated behavior, as shown in Fig. 3d, while the temperature dependence of $G_{\text{PF}}$ suggests a robust charge gap. This strongly suggests that these doping-induced excitons are intrinsic features of the ground state rather than thermal excitations.
Notably, tuning the total filling fraction, $\nu_{\text{total}}$, results in a region with a zero drag ratio, outlined by blue dashed lines and highlighted by the blue-shaded area in the middle and right panels of Fig. 3b, respectively. Fig. 3e plots the transport responses as $\nu_{1}$ is detuned from the Laughlin state. The onset of $G_{\text{PF}}$ indicates that the charge gap is suppressed by the presence of quasiparticles in layer 1 (black solid line in Fig. 3e). However, the absence of quasiholes in layer 2 prevents the formation of exciton pairs, thereby suppressing the counterflow conductance, $G_{\text{CF}}$ (blue solid line in Fig. 3e). This results in a non-zero $G_{\text{PF}}$ and a zero $G_{\text{CF}}$, confirming that exciton formation requires doping with varying $\Delta\nu$ while maintaining a constant $\nu_{\text{total}}$.
Having established quasiparticle pairing in the Laughlin state, we now explore other neutral modes by examining transport properties along the equal density line, which is defined by $\nu_{1}=\nu_{2}$ and depicted by the blue horizontal line in Fig. 4a. Along this line, transport measurement uncovers three distinct types of neutral modes in the form of bilayer excitons, as shown in Fig. 4b-d, with their corresponding locations marked by green, blue, and red circles in Fig. 4a. At $\nu_{1}=\nu_{2}=1/2$ (green circle), the perfect drag response, accompanied by vanishing $G_{\text{PF}}$, points to an excitonic state described by the (111) wavefunction  Nandi2012exciton ; Eisenstein2014 ; Zeng2023solid ; Zhang2024fractionalexciton ; Li2017superfluid ; Liu2017superfluid ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . These excitons consist of components with integer charge, which combine to form composite bosons (see inset in Fig. 4b, framed by a green square). As $\nu_{1}$ and $\nu_{2}$ deviate from half filling, a series of incompressible states–evidenced by vanishing $G_{\text{PF}}$–emerge at positions expected from the Jain sequence, defined by $\nu_{1}^{\ast}=\nu_{2}^{\ast}=N$, with $N\in\mathbb{Z}$.
This sequence of Jain states exhibit two interconnected hierarchical behaviors. First, the magnitude of the charge gap, reflected by the width of the insulating features in $G_{\text{PF}}$, is enhanced at lower effective filling. Second, quasiparticle pairing is suppressed at $\nu_{1}=\nu_{2}=1/3$ due to the large charge gap, but re-emerges as the charge gap decreases at higher effective filling. This is demonstrated by the perfect drag response and non-zero counterflow conductance at $\nu_{1}=\nu_{2}=2/5$, $3/7$, and $4/9$. As shown in the insets of Fig. 4b (framed by blue squares), these excitons have fractionally charged constituents that obey anyonic statistics, leading the neutral modes to exhibit anyonic behavior as well.
Most remarkably, transport measurement unveils an unexpected neutral mode at $\nu_{1}=\nu_{2}=3/8$, characterized by vanishing $G_{\text{PF}}$, enhanced counterflow conductance, $G_{\text{CF}}$, and a peak in the drag ratio approaching unity. Two observations suggest that this state is linked to the Jain sequences of FQHE where composite fermions (CFs) are constructed with two and four intralayer flux attachments, but notably without interlayer flux attachments  Li2019pairing ; Zhang2024fractionalexciton .
First, the state at $\nu_{1}=\nu_{2}=3/8$ is surrounded by these Jain sequences of FQHE states, which appears at effective integer filling of CFs. These Jain states, marked by blue, black, and gray circles in Fig. 4a, are identified by the vanishing $G_{\text{PF}}$ (dark blue) in Fig. 4f. Second, the FQHE state at $\nu_{1}=\nu_{2}=3/8$ exhibits a $B$-induced transition similar to transitions observed at $\nu_{1}=\nu_{2}=1/3$, $2/5$, and $3/7$, as marked by black horizontal lines in Fig. 4e. The low-$B$ states are described by two-component wavefunctions  Li2019pairing ; Zhang2024fractionalexciton , whereas the high-$B$ states are one-component in nature, which fall into the conventional Jain sequence.
Despite its connection with the Jain sequence, the state at $\nu_{1}=\nu_{2}=3/8$ appears at a half-filled $\Lambda$ level, suggesting an exotic origin driven by residual interactions between CFs  Pan.03 ; xia2004electron ; pan2015fractional ; bellani2010optical ; samkharadze2015observation . The emergence of this state may follow a similar mechanism as the celebrated $\nu=5/2$ state. In this case, electrons in the lowest LL capture two-flux quanta and populate $\Lambda$ levels in response to the residual magnetic field. When the second $\Lambda$ level is half-filled, the two-flux composite fermions capture two additional flux quanta, undergo Cooper instability, and condensed into a $p-$wave paired state. This scenario is supported by previous numerical simulations, identifying a relative to the Bonderson-Slingerland (BS) wavefunction as the stable ground state at $\nu=3/8$ in the lowest LL of a single 2D confinement  scarola2002possible ; Mukherjee2012Pfaffian ; mukherjee2014possible ; Bonderson2008 . This describes a non-Abelian state similar to the anti-Pfaffian state at $\nu=5/2$.
Intriguingly, previous studies on ultra-high mobility GaAs samples have reported the FQHE at $\nu=3/8$, originating from residual composite fermion interaction  wang2022even ; wang2023next ; wang2023fractional . However, this exotic state has been notably absent in prior experimental studies involving monolayer graphene samples with similar doubly-encapsulated geometry  Zibrov2018even ; Zeng2019 ; Chen2019 ; Polshyn2018corbino . It should be noted that LL mixing plays a significant role in facilitating interaction between composite fermions, creating favorable condition for pairing at half-filled $\Lambda$ levels  zhao2023composite ; wang2023next . Indeed, these unconventional states in GaAs were observed in hole-doped LL filling rather than electron-doped, presumably because the larger effective mass of holes, and consequently larger LL mixing, enhances interactions between CFs wang2022even ; wang2023next ; wang2023fractional . In quantum Hall graphene bilayers, we expect that the unique Coulomb environment reduces short-range repulsion in the CF sea, making it more favorable for CFs to form Cooper pairs. At large magnetic field, the absence of interlayer correlations restricts Cooper pairing to occur within each individual layer. As a result, the bilayer system offers a favorable platform to stabilize novel composite fermion states driven by intra-layer correlations, which are distinct from those observed in monolayer graphene samples.
At $\nu=3/8$, quasiparticle excitations in the sample bulk likely obeys non-Abelian quantum statistics, which arises from the non-trivial fusion rules between the distinct topological sectors. In a quantum Hall bilayer, we expect that quasiparticles at $\nu_{1}=\nu_{2}=3/8$ belong to the same topological sector. This results in a bilayer state with charge-neutral excitons that follow non-Abelian statistics, a highly unusual construction (see Methods section for more detailed discussion on the likely non-Abelian order of the bilayer state).
In summary, our findings reveal charge-neutral anyons formed through the pairing of quasiparticles and quasiholes. These novel anyons, emerging from the well-studied Jain sequence of FQHE states, highlight the charge-neutral channel as an unexplored pathway that has remained largely hidden until this work. Neutral anyons represent a new class of quantum particles distinct from charged particles. Their charge neutrality leads to a unique form of correlation beyond conventional Coulomb interaction, especially in the context of bilayer excitons. The flow dynamics of neutral anyons can be directly explored using counterflow transport, while their population can be uniquely controlled through layer-asymmetric field-effect doping. Together, these properties and tunability open a new dimension for constructing and engineering exotic quantum phases of matter, which will certainly inspire extensive future research interests.
acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. N.J.Z., R.Q.N., and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. N.J.Z. acknowledges partial support from the Jun-Qi fellowship. R.Q.N. and J.I.A.L. acknowledges partial support from the National Science Foundation EPSCoR Program under NSF Award OIA-2327206. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Alfred P. Sloan Foundation (G-2024-22395). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-1157490 and the State of Florida. N.B. and D.E.F. were supported in part by the National Science Foundation under Grant No. DMR-2204635. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan.
Data availability
The data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
Author contribution
N.J.Z. and J.I.A.L. conceived the project. N.J.Z., R.Q.N., and X.L. fabricated the device. N.J.Z. and R.Q.N. performed the measurement. N.B. and D.E.F. provided theoretical inputs. K.W. and T.T. provided the material. N.J.Z., R.Q.N., N.B., D.E.F., and J.I.A.L wrote the manuscript together.
Competing financial interests
The authors declare no competing financial interests.
Supplementary Text
II Supplementary Text
II.1 Fractional charge and statistics of excitons belonging to ${}^{2}_{0}$ family
In this section we find the fractional charge and self-statistical phase for the Abelian interlayer excitons considered in the main text. We limit to the case when the effective filling $\nu_{i}^{*}\in\mathbb{Z}$ for each layer $i=1,2$. Since there are no interlayer fluxes, the bilayer ${\bf K}$-matrix is block diagonal ${\bf K}=\text{diag}({\bf K}_{1},{\bf K}_{2})$ with each block ${\bf K}_{i}$ being the monolayer ${\bf K}$-matrix along with their corresponding charge vector ${\bf t}_{i}$. Working in the symmetric representation ${\bf t}=({\bf t}_{1}\oplus{\bf t}_{2})^{T}=(1,\dots,1)$, each monolayer ${\bf K}$-matrix takes a simple form,
${\bf K}_{i}=\sigma_{i}{\bf I}+2{\bf J}~{}~{}~{}~{}\Leftrightarrow~{}~{}~{}~{}{\bf K}^{-1}_{i}=\sigma_{i}{\bf I}-\frac{2\sigma_{i}}{2\dim{\bf K}_{i}+\sigma_{i}}{\bf J}~{}~{}~{}~{}~{}~{}~{}\text{for}~{}~{}i=1,2,$ (S1)
where $\sigma_{i}=\pm 1$. The matrices ${\bf I}$ and ${\bf J}$ are identity and a constant matrix of ones respectively. Since $\dim{\bf K}_{i}$ represent the number of $U(1)$ gauge fields in the low energy Chern-Simons description, we can set $\dim{\bf K}_{i}=|\nu_{i}^{*}|$. Note that this identification only works for the Abelian states when $\nu_{i}^{*}\in\mathbb{Z}$ which we focus on in this section. Similarly, $\sigma_{i}$ is chosen depending upon the ${}^{2}_{0}$ particle-/hole-doped $\Lambda$ levels so we identify $\sigma_{i}=\text{sign}(\nu_{i}^{*})$.
Next, we compute the fractional charge for an arbitrary excitation labeled by ${\bf l}_{i}$ in the $i$th layer,
$Q_{i}=-e~{}{\bf t}^{T}_{i}{\bf K}_{i}^{-1}{\bf l}_{i}$ $=-e\sum_{n,m}^{|\nu_{i}^{*}|}\left(\sigma_{i}\delta_{nm}-\frac{2\sigma_{i}}{2|\nu_{i}^{*}|+\sigma_{i}}J_{nm}\right)\left[{\bf l}_{i}\right]_{m}=-e\left(\sigma_{i}-\frac{2\sigma_{i}}{2|\nu_{i}^{*}|+\sigma_{i}}|\nu_{i}|^{*}\right)\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$ (S2)
$=-e\frac{\ell_{i}}{\sigma_{i}\left(2\nu_{i}^{*}+1\right)},$ (S3)
where in the last equality we used $\sigma_{i}=\text{sign}(\nu_{i}^{*})$ and defined the integer $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$. Since the vector ${\bf l}_{i}$ has integer components, the minimal charge excitation in each layer is $Q_{\text{min.},i}=\pm e/|2\nu_{i}^{*}+1|$ which is simply the fundamental charge in the Jain states. Next, we find the minimal exciton charge for the bilayer system via the condition $Q_{1}+Q_{2}=0$ such that $Q_{1,2}\neq 0$. The minimal integer solutions of $\ell_{1},\ell_{2}\in\mathbb{Z}$ are,
$\frac{\ell_{1}}{|2\nu_{1}^{*}+1|}+\frac{\ell_{2}}{|2\nu_{2}^{*}+1|}=0~{}~{}~{}~{}\Rightarrow~{}~{}~{}~{}\ell_{i}=\frac{(-1)^{i}~{}|2\nu_{i}^{*}+1|}{\gcd(|2\nu_{i}^{*}+1|,|2\nu_{2}^{*}+1|)}~{}~{}~{}i=1,2,$ (S4)
where the integer valued function $\text{gcd}(n_{1},n_{2})$ computes the greatest common divisor of the integers $n_{1}$ and $n_{2}$. We also use $\text{sign}(\nu_{i}^{*})=\text{sign}(2\nu_{i}^{*}+1)$ since $|\nu_{i}^{*}|\geq 1$. This gives the minimal exciton charge,
$Q_{\text{exciton}}=\pm\frac{e}{\gcd(|2\nu_{i}^{*}+1|,|2\nu_{2}^{*}+1|)}.$ (S5)
The minimal solutions $\ell_{i}$ can also be used to compute the exciton statistics. Before we compute the exciton statistics, we notice that due to the definition $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$, there is an ambiguity in defining the excitation corresponding to the integer vector ${\bf l}_{i}$. This ambiguity suggests that there are distinct solutions of ${\bf l}$ that all correspond to the same exciton charge, or for that matter, a generic quasiparticle charge. However, to make sure all these distinct solutions describe the same particle, we need to check if all such excitations result in the same mutual and self statistical phases modulo $2\pi$. We proved that the distinct solutions indeed result in the same statistical phases in our previous report (see Ref. Zhang2024fractionalexciton ) for a general bilayer ${}^{a}_{b}$ family. This guarantees that for any choice of ${\bf l}_{i}$ corresponding to the solutions obtained $\ell_{i}$ indeed describes the same minimal exciton as with any other choice of ${\bf l}_{i}$ preserving $\ell_{i}$. We thus set $\left[{\bf l}_{1}\right]_{1}=\ell_{1}$ and $\left[{\bf l}_{2}\right]_{1}=\ell_{2}$, and all the remaining elements are set to zero $[{\bf l}_{i}]_{n}=0$ $\forall~{}n\in[2,|\nu_{i}^{*}|]$. The exciton self-statistics is computed as,
$\theta_{s}$ $=\pi\sum_{i=1}^{2}{\bf l}_{i}^{T}{\bf K}_{i}^{-1}{\bf l}_{i}\mod 2\pi=\pi\left[{\bf K}_{1}^{-1}\right]_{11}\ell_{1}^{2}+\pi\left[{\bf K}_{2}^{-1}\right]_{11}\ell_{2}^{2}\mod 2\pi$ (S6)
$=\frac{\pi}{\gcd(|2\nu_{i}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}}\sum_{i=1}^{2}\left[\sigma_{i}-\frac{2}{(2\nu_{i}^{*}+1)}\right](2\nu_{i}^{*}+1)^{2}\mod 2\pi$ (S7)
$=\pi\sum_{i=1}^{2}\frac{\left(|2\nu_{i}^{*}+1|-2\right)\left(2\nu_{i}^{*}+1\right)}{\text{gcd}(|2\nu_{1}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}}\mod 2\pi$ (S8)
which shows that the excitons, in general, follow anyonic statistics rather than the usually anticipated bosonic character. Interestingly, we notice that since $\nu_{i}^{*}\in\mathbb{Z}$, the sum $2x\equiv\sum_{i=1}^{2}\left(|2\nu_{i}^{*}+1|-2\right)\left(2\nu_{i}^{*}+1\right)\in 2\mathbb{Z}$ for some integer $x\in\mathbb{Z}$. Similarly, the denominator $12y+1\equiv\text{gcd}(|2\nu_{1}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}\in 2\mathbb{Z}+$ 1 for some $y\in\mathbb{Z}$. Now since the following equation has no integer solutions,
$\frac{x}{2y+1}\mod 1=\frac{1}{2}~{}~{}~{}~{}\Rightarrow~{}~{}~{}~{}\text{no solution for $x,y\in\mathbb{Z}$},$ (S9)
the excitons in the above construction are never fermionic for any combination of monolayer Jain states $\nu_{i}^{*}\in\mathbb{Z}$.