### v2

# Bilayer Excitons in the Laughlin Fractional Quantum Hall State

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: Email: jia__\__li@brown.edu
Bilayer Excitons in the Laughlin Fractional Quantum Hall State
Ron Q. Nguyen1    Naiyuan J. Zhang1    Navketan Khurana-Batra1,2    Xiaoxue Liu3    Kenji Watanabe4    Takashi Taniguchi5    D. E. Feldman1,2    J.I.A. Li1 1Department of Physics, Brown University, Providence, Rhode Island 02912, USA 2Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA 3Tsung-Dao Lee Institute and School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China 4Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 5Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(July 5, 2025)
The Laughlin state embodies a universal class of fractional quantum Hall effects arising in two-dimensional electron systems subjected to strong perpendicular magnetic fields. Conventionally described by a single-component wavefunction, the Laughlin state features fractionally charged quasiparticles arising from correlations within one electron species. Here, we explore a novel physical situation by introducing inter-species Coulomb coupling between two intra-species Laughlin states in a quantum Hall graphene bilayer structure. Although quasiparticle excitations typically exhibit charge gaps of tens of Kelvin, we observe that this energy scale is significantly lowered through interlayer excitonic pairing between quasiparticles and quasiholes. Identified via transport measurements, these excitons belong to an unprecedented category of charge-neutral anyons, opening a new avenue for investigating exotic quantum statistics and phases of matter.
Figure 1: Exciton pairing between a pair of Laughlin states. (a) Schematic of the Laughlin state in a monolayer, stabilized by intralayer Coulomb interaction. At $1/3$ LL filling, quasiparticle excitations of the Laughlin state carry $1/3$ of electronic charge. (b) Schematic of two Laughlin states in a quantum Hall bilayer, each stabilized by intralayer Coulomb interaction. In this situation, quasiparticles and quasiholes interact through weak interlayer correlation. (c) Schematic diagram showing FQHE within the CF model, where the attachment of two flux quanta is included in the CF construction  Jain.03 . Blue and red solid lines denote integer values for $\nu^{\ast}_{1}$ and $\nu^{\ast}_{2}$, respectively  Li2019pairing ; Liu2019interlayer . At the intersection of these lines, both layers simultaneously develop a robust charge gap, leading to an incompressible state across the quantum Hall bilayer. (d-e) Transport response along $\nu_{1}^{\ast}=1$, at (d) $B=12$ T and (e) $B=28$ T. The top, middle, and bottom panels show PF conductance $G_{\text{PF}}$, counterflow conductance $G_{\text{counter}}$, and drag ratio $I_{\text{drag}}/I_{\text{drive}}$, respectively. Inset of each panel shows schematic measurement configurations in the parallel flow, counterflow, and drag geometries. Blue, green, and red circles mark different ground state orders.
Figure 2: Pairing between quasiparticle excitations. (a,b) Schematic diagrams illustrating (a) thermally excited quasiparticles and quasiholes across the charge gap of the Laughlin state, and (b) suppressed quasiparticle excitation when the charge gap greatly exceeds $k_{B}T$. (c,d) Temperature dependence of (c) drive and drag currents, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$, in a drag measurement, and (d) the drag ratio $I_{\textrm{drag}}/I_{\textrm{drive}}$. Measurements are performed at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. The gray shaded region indicates the low-temperature regime where both currents vanish. The red shaded region marks the onset of simultaneous $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$, yielding a perfect drag ratio. At higher temperatures, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ bifurcate, and the drag ratio deviates from unity. (e) Magnetic field dependence of the counterflow conductance $G_{\text{counter}}$ (log scale) measured at $\nu_{1}=\nu_{2}=1/3$ (black circles), $2/5$ (blue circles), and $3/7$ (red circles).
Quantum statistics form the cornerstone for understanding diverse quantum phases of matter in condensed matter physics. Fermionic statistics fundamentally describe electronic systems at low temperatures, underpinning the physics of Fermi liquids Luttinger1960fermi ; Nozieres2018quantumliquids and the quantum Hall effect Halperin2020fractional . When fermions pair, they form composite particles which are bosons, enabling the emergence of low-temperature ground states characterized by Bose–Einstein condensation. Superconductivity, for instance, results from electrons near Fermi surface undergoing Cooper pairing instabilities Cooper2011BCS . Alternatively, excitonic superfluids represent a distinct class of bosonic condensate, resulting from Coulomb-driven electron-hole pairing, establishing an intriguing phenomenon that is actively pursued in condensed matter research Blatt1962exciton ; Lozovik1975exciton ; Lozovik1976exciton .
While excitons are accessible in a wide range of materials, quantum Hall bilayers represent a paradigmatic solid-state platform for stabilizing excitonic condensates. In these bilayer structures, charge carriers reside in two closely spaced but electrically isolated two-dimensional layers. Coulomb attraction across the layers drives electron-hole pairing, stabilizing various emergent excitonic phases extensively studied both experimentally and theoretically Nandi2012exciton ; Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Li2019pairing ; Liu2022crossover ; Eisenstein2004BEC ; Kellogg2005 ; Zeng2023solid ; Zhang2025fractionalexciton ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . Among the most extensively examined phases, an excitonic condensate emerges at a total Landau level (LL) filling factor of one, accurately described by the (111) wavefunction Halperin1983 ; Wen1992 , representing an excitonic analogue of the integer quantum Hall effect state.
More recently, it was discovered that interlayer excitons can coexist with fractional quantum Hall states, indicating that excitonic phenomena are more ubiquitous within the quantum Hall regime than previously recognized. In the regime with strong interlayer correlation, fractional quantum Hall states can be described by two-component wavefunctions, notably the (333) state Halperin1983 ; Wen1992 , serving as the fractional analogue to the (111) state. Here, interlayer excitons naturally emerge as essential constituents of the ground-state construction Zhang2025fractionalexciton .
In this work, we report a new pathway for creating interlayer excitons by introducing Coulomb coupling between two Laughlin states. While the Laughlin state is traditionally regarded as a single-component state, we show that Coulomb coupling between two such states generates interlayer excitons at energies substantially lower than the charge gap. As shown in Fig. 1a, quasiparticle excitation of the Laughlin state carries $1/3$ of an electron charge and obeys anyonic quantum statistics Tsui1982FQHE ; halperin1984 ; Arovas1984fractional ; review-FH ; Picciotto1997 ; Saminadayar1997 ; Reznikov1999 ; Bartolomei2020fractional ; Nakamura2020Anyon ; texp1 ; texp2 ; Willett2023interference . We propose that interlayer excitons of two Laughlin states results from pairing between fractionally charged quasiparticles and quasiholes (Fig. 1b).
The phase space of a quantum Hall bilayer is defined by the Landau level fillings of each graphene layer, $\nu_{1}$ and $\nu_{2}$, which can be independently controlled by adjusting the voltage bias on the top and bottom gate electrodes. Figure 1c presents a schematic phase diagram, where the axes are redefined by the total filling $\nu_{\text{total}}=\nu_{1}+\nu_{2}$ and the layer imbalance $\Delta\nu=\nu_{1}-\nu_{2}$. We focus on the weak-coupling regime dominated by the conventional Jain sequence of FQHE states. Here, red and blue solid lines indicate the expected locations of single-component FQHE states in layers 1 and 2, respectively. Black circles mark a representative set of intersections, where layer 1 is tuned to a Laughlin state at $\nu_{1}=1/3$ and layer 2 is swept across a sequence of Jain states characterized by the effective composite fermion (CF) filling $\nu_{2}^{\ast}=\nu_{2}/(1-2\nu_{2})=N$ with $N\in\mathbb{Z}$ Jain.03 ; Eisenstein1990FQHE . These CFs are constructed using intralayer flux attachment only, distinguishing them from previous reports of two-component FQHE Li2019pairing ; Zhang2025fractionalexciton . In the following, we use $\nu_{i}^{\ast}$ to denote the CF effective filling in layer $i=1,2$.
Across the Jain sequence highlighted in Fig. 1c, the presence of interlayer excitons is revealed through their distinctive transport responses. We employ three distinct electrical transport configurations in bilayer samples patterned into an edgeless Corbino geometry (see Methods for details) Nandi2012exciton ; Eisenstein2014 ; Li2019pairing ; Zeng2023solid ; Zhang2025fractionalexciton . At $B=12$ T, the parallel-flow (PF) geometry exhibits vanishing bulk conductance at $\nu_{2}^{\ast}=N$ (top panel of Fig. 1d), confirming robust charge gaps associated with Jain states. Remarkably, these insulating states in the PF geometry exhibit non-zero conductance in the counterflow configuration (middle panel of Fig. 1d), coinciding with perfect drag responses (bottom panel of Fig. 1d). When a current bias is applied to the drive layer (layer 1), a counter-propagating current of equal magnitude is detected in the drag layer (layer 2), resulting in a unity drag ratio $I_{\text{drag}}/I_{\text{drive}}=1$. This combination of transport signatures is in excellent agreement with prior observations of the (111) state, offering unambiguous evidence for the existence of interlayer excitons across the Jain sequence.
Interestingly, signatures of exciton pairing between a pair of Laughlin state vanishes at $B=28$ T. As shown in Fig. 1e, vanishing conductance is observed in both the PF and counterflow geometries at $\nu_{1}=\nu_{2}=1/3$ (top and middle panels of Fig. 1e). In the drag geometry, both the drive and drag currents approach zero, leading to an ill-defined drag ratio, which is masked with the gray-shaded region in the bottom panel of Fig. 1e.
The Jain sequence of FQHE states, characterized by its hierarchical behavior shown in Fig. 1d-e, is described by well-defined wavefunctions localized within each layer. The absence of interlayer correlation in the ground-state wavefunction raises an intriguing conundrum regarding the origin of interlayer excitons.
The magnetic field dependence of the transport response, as illustrated in Fig. 1d-e, provides an important clue. We propose that exciton pairing occurs between a negatively charged quasiparticle in one layer and a positively charged quasihole in the other. As shown in the inset of Fig. 2a, such quasiparticles and quasiholes are thermally generated across the charge gap. Consequently, their population is governed by the temperature relative to the charge gap. In systems with a large charge gap, such as the situation at $B=28$ T, these excitons are suppressed at low temperatures but are expected to re-emerge as the gap diminishes or as the temperature increases.
This hypothesis is validated by the temperature dependence of the drag response. At $B=28$ T and $\nu_{1}=\nu_{2}=1/3$, drag measurement produces zero $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ in the low-temperature region, as shown in Fig. 2c. This is consistent with the absence of excitons, and both layers behave like insulators in the bulk. With increasing temperature, quasiparticle excitations are generated across the charge gap, evidenced by non-zero current in both drive and drag layers. Remarkably, the onset of current across two layers recovers a perfect drag ratio, as shown in the red shaded area in Fig. 2d. This observation offers strong support for exciton pairing from thermally excited quasiparticle and quasihole. Upon further increasing $T$, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ bifurcate and the drag ratio start to deviate from $1$, indicating that exciton pairing is suppressed at high temperature.
Since the transport response in the parallel flow configuration does not couple to interlayer excitons, the activated behavior in $G_{\text{PF}}$ offers a direct measurement of the underlying charge gap. As illustrated by Fig. S1, the charge gap associated with the Laughlin state at $B=28$T is around $\Delta=40.5$ K, substantially higher than the temperature of current onset in the drag measurement. This discrepancy suggests that exciton pairing lowers the energy cost for creating quasiparticle excitations.
Moreover, the Jain states follow a hallmark hierarchical behavior. As the effective magnetic field decreases at larger effective filling, the charge gap is reduced, making it more preferable to generate quasiparticles and quasiholes. This has a direct impact on the transport response in the counterflow channel. As shown in the bottom panel of Fig. 1e, while the Laughlin state exhibits zero counterflow conductance at $B=28$ T, a perfect drag response is recovered as layer 2 is tuned to Jain states at larger effective fillings.
Along the same vein, Fig. 2e plots counterflow conductance $G_{\text{counter}}$ as a function of $B$ measured at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=1$, $2$, and $3$. At $\nu_{1}^{\ast}=\nu_{2}^{\ast}=1$, exciton population disappears after $B=15$ T, evidenced by vanishing $G_{\text{counter}}$. On the other hand, exciton pairing persists up to $B$ = 30 T at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=2$ and even higher $B$ at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=3$. This comparison establishes a general behavior: at each filling, exciton pairing is suppressed as the charge gap is enhanced with increasing $B$, while at each $B$, excitons become more stable as the charge gap is suppressed at larger effective filling. Combined, our observations strongly support the proposed pairing mechanism between quasiparticles.
Figure 3: Doping neutral anyons into the Laughlin state. (a) Controlling the population of neutral anyons through layer-asymmetric doping, which is achieved by varying Landau level fillings of layer 1 and 2 by ±$\pm\delta$. This generates equal number of quasiparticles and quasiholes across two graphene layers. (b) Parallel flow conductance $G_{\text{PF}}$ (left), drag ratio $I_{\text{drag}}/I_{\text{drive}}$ (middle), and schematic diagram (right) summarizing the transport response across the $\nu_{\text{total}}$-$\Delta\nu$ map around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. Near $\Delta\nu$=0, $I_{\text{drag}}$ and $I_{\text{drive}}$ both vanish, resulting in ill-defined drag ratio masked by gray-shaded areas. Black arrows in the right panel mark the extend of the robust charge gap along the direction of $\Delta\nu$ and $\nu_{\text{total}}$. (c) $G_{\text{PF}}$, $G_{\text{counter}}$ (top) and $I_{\text{drag}}/I_{\text{drive}}$ (bottom) as a function of $\Delta\nu$ measured at $\nu_{\text{total}}=2/3$. (d) Arrhenius plot of $G_{\text{PF}}$ (black) and $G_{\text{counter}}$ (blue) measured at $\Delta\nu=0.06$. (e) $G_{\text{PF}}$, $G_{\text{counter}}$ as a function of varying $\nu_{1}$ while $\nu_{2}$ is fixed at $1/3$.
Figure 4: Non-Abelian neutral mode at $\nu_{1}=\nu_{2}=3/8$. (a) A detailed view of the $\nu_{\text{total}}$-$\Delta\nu$ phase space, where black and gray lines indicate CF states with two and four flux attachments, respectively. (b) Parallel flow conductance $G_{\text{PF}}$, (c) counterflow conductance $G_{\text{counter}}$, and (d) drag ratio $I_{\text{drag}}/I_{\text{drive}}$ as a function of varying $\nu_{\text{total}}$ measured at $B=31$ T. Vertical arrows highlight different types of charge-neutral anyons emerging at locations marked by red, blue, and green circles in panel (a). (e) $G_{\text{PF}}$ as a function of $\nu_{\text{total}}$ and $B$, revealing a series of $B$-induced transitions in the FQHE states, marked by horizontal black solid lines. (f) $G_{\text{PF}}$ measured across a portion of the phase space encircled by the red dashed line in panel (a). Vanishing $G_{\text{PF}}$ defines incompressible states marked by black and gray circles in panel (a).
In the above mentioned scenario, quasiparticles, quasiholes, and the resulting excitons are all thermal excitations. In the following, we show that such excitons can also be generated as a feature of the low-temperature ground state, which is realized through independently control field-effect doping across two graphene layers.
As illustrated in Fig. 3a, an equal population of quasiparticles and quasiholes can be generated across the two graphene layers by doping the bilayer to $\nu_{1}=1/3+\delta$ and $\nu_{2}=1/3-\delta$. This defines a vertical line in phase space at constant total filling fraction $\nu_{\text{total}}=2/3$, where $\delta$ is directly linked to layer imbalance $\Delta\nu$.
The presence of doping-induced excitons is demonstrated by the transport properties around $\nu_{1}=\nu_{2}=1/3$. Vanishing parallel flow conductance with $G_{\text{PF}}=0$ highlights a region with a fully developed charge gap, represented by dark blue in the left panel of Fig. 3b. While this region spans a range of $0.20$ in the direction of $\Delta\nu$, it covers a much narrower range of only $0.02$ in the direction of $\nu_{\text{total}}$. This directional dependence aligns well with the unique doping requirement for exciton formation. Tuning along other directions, such as varying $\nu_{\text{total}}$ while holding $\Delta\nu$ fixed, leads to mismatched quasiparticle and quasihole populations, resulting in unpaired charges and a lack of drag response.
Within the boundary of vanishing $G_{\text{PF}}$, drag measurements reveal the emergence of exciton pairing. Near $\Delta\nu$ $=0$, the bilayer sample exhibits vanishing conductance in both parallel flow and counterflow measurements, as illustrated in Fig. 3c. As both drive and drag currents are zero, the drag ratio is ill-defined, which is indicated by the gray-shaded area in Fig. 3b and c. As $\Delta\nu$ deviates from zero, the counterflow conductance rises from zero while the parallel flow conductance remains vanishingly small. This leads to a regime characterized by a perfect drag response, represented by dark red in the middle and right panels of panel of Fig. 3b and indicated in the red-shaded area of Fig. 3c. These transport responses point to excitonic pairing, driven by the doping-induced population of quasiparticles and quasiholes.
In the presence of doping-induced excitons, the bulk conductance in the counterflow geometries no longer exhibits activated behavior, as shown in Fig. 3d, while the temperature dependence of $G_{\text{PF}}$ suggests a robust charge gap. This strongly suggests that these doping-induced excitons are intrinsic features of the ground state rather than thermal excitations.
Notably, tuning $\nu_{\text{total}}$ at $\Delta\nu$$=0$ results in a region with a zero drag ratio, outlined by blue dashed lines and highlighted by the blue-shaded area in the middle and right panels of Fig. 3b, respectively. Fig. 3e plots the transport responses as $\nu_{1}$ is detuned from the Laughlin state. The onset of $G_{\text{PF}}$ indicates that the charge gap is suppressed by the presence of quasiparticles in layer 1 (black solid line in Fig. 3e). However, the absence of quasiholes in layer 2 prevents the formation of exciton pairs, thereby suppressing the counterflow conductance, $G_{\text{counter}}$ (blue solid line in Fig. 3e). This results in a non-zero $G_{\text{PF}}$ and a zero $G_{\text{counter}}$, confirming that exciton formation requires doping with varying $\Delta\nu$ while maintaining a constant $\nu_{\text{total}}$.
Having established quasiparticle pairing in the Laughlin state, we now expand our investigation to other types of excitons along the equal density line, defined by $\nu_{1}=\nu_{2}$ and depicted as the horizontal blue line in Fig. 4a. Along this line, transport measurements reveal three distinct types of excitonic states, with their respective positions marked by green, blue, and red circles in Fig. 4a. At $\nu_{1}=\nu_{2}=1/2$ (green circle), the perfect drag response accompanied by vanishing $G_{\text{PF}}$ indicates an excitonic state described by the (111) wavefunction Nandi2012exciton ; Eisenstein2014 ; Zeng2023solid ; Zhang2025fractionalexciton ; Li2017superfluid ; Liu2017superfluid ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . These excitons are composed of integer-charged constituents that combine to form composite bosons (see inset in Fig. 4b, green frame).
As $\nu_{1}$ and $\nu_{2}$ deviate from half-filling, a series of FQHE states emerge along the Jain sequence, as evidenced by vanishing $G_{\text{PF}}$ at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=2$, 3, 4, and 5 (Fig. 4b). At these fillings, we observe a perfect drag ratio and non-zero counterflow conductance, providing strong support for the presence of interlayer excitons. We argue that the pairing mechanism at these fillings is analogous to that at $\nu_{1}=\nu_{2}=1/3$, where excitons form via interlayer pairing of quasiparticles and quasiholes.
Notably, the interlayer excitons observed along the Jain sequence should be distinguished from those described by the (111) state at $\nu_{1}=\nu_{2}=1/2$ Eisenstein2014 ; Li2017superfluid ; Liu2017superfluid ; Liu2022crossover , as well as those associated with two-component fractional quantum Hall effect Zhang2025fractionalexciton . Ground states described by two-component wavefunctions, such as the (111) and (333) states Halperin1983 ; Wen1992 , arise in regimes where strong interlayer correlation plays a critical role in defining the FQHE phase. In contrast, the regime explored in this work is characterized by weak interlayer coupling, where FQHE states are stabilized primarily by intralayer Coulomb interactions.
The regimes of strong and weak interlayer correlation can be identified based on the hierarchical behavior of the FQHE sequence Jain.03 ; Li2019pairing . The boundary between these two regimes are revealed in Figure 4e, which plots $G_{\text{PF}}$ as a function of LL filling and $B$, measured along the equal density line. The regime of strong interlayer correlation occurs at $B<10$ T, where the FQHE sequence aligns with composite fermions incorporating interlayer flux attachment Li2019pairing ; Zhang2025fractionalexciton . In contrast, at $B>15$ T, the FQHE sequence exhibits the characteristic hierarchical structure of the Jain sequence. These two regimes are separated by a series of $B$-induced transitions at $\nu_{1}=\nu_{2}=1/3$, $2/5$, and $3/7$, marked by black horizontal lines in Fig. 4e, where the charge gap is weakened due to a transition between two-component and one-component FQHE ground states.
Amongst the Jain sequence of FQHE states in the regime of weak interlayer correlation, transport measurements unveil an unexpected FQHE state at $\nu_{1}=\nu_{2}=3/8$. This state emerges at $B=12$ T and persists beyond $B=31$ T (Fig. 4b-e), offering strong support that the underlying charge order is single-component in nature. However, $\nu_{1}=\nu_{2}=3/8$ lies outside the Jain sequence of FQHE, indicating an exotic type of order.
In the following, we propose a plausible description of relevant physics at $\nu_{1}=\nu_{2}=3/8$.
We first discuss the ground state wavefunction in each graphene layer, which defines the nature of the charge gap. Since this filling corresponds to half-filled $\Lambda$ level of component fermions  Pan.03 ; xia2004electron ; pan2015fractional ; bellani2010optical ; samkharadze2015observation , the emergence of a robust FQHE state could follow a similar mechanism as the celebrated $\nu=5/2$ state. In this case, electrons in the lowest LL capture two-flux quanta and populate $\Lambda$ levels in response to the residual magnetic field. When the second $\Lambda$ level is half-filled, the two-flux composite fermions capture two additional flux quanta, undergo Cooper instability, and condensed into a $p-$wave paired state. This scenario is supported by previous numerical simulations, identifying a relative to the Bonderson-Slingerland (BS) wavefunction as the stable ground state at $\nu=3/8$ in the lowest LL of a single 2D confinement  scarola2002possible ; Mukherjee2012Pfaffian ; mukherjee2014possible ; Bonderson2008 . This describes a non-Abelian state similar to the Pfaffian state at $\nu=5/2$  Halperin2020fractional . In this scenario, quasiparticle excitations in the sample bulk likely obeys non-Abelian quantum statistics, which arises from the non-trivial fusion rules between the distinct topological sectors.
Most interestingly, the transport response at $\nu_{1}=\nu_{2}=3/8$ offers clear indication for interlayer exciton pairing. Besides vanishing $G_{\text{PF}}$, this state exhibits an enhanced counterflow conductance, as well as a peak in the drag ratio approaching unity. Given the robust charge gap and the single-component nature of the ground state, we argue that exciton pairing occurs between quasiparticle and quasihole excitations. Notably, if the FQHE state at $\nu_{1}=\nu_{2}=3/8$ is non-Abelian in nature, the resulting exciton will follow non-Abelian statistics as well, giving rise to a highly unusual construction (see Methods section for more detailed discussion on the likely non-Abelian order of the bilayer state).
Previous studies on ultra-high mobility GaAs samples have reported the FQHE at $\nu=3/8$, originating from residual composite fermion interaction  wang2022even ; wang2023next ; wang2023fractional . However, this exotic state has been notably absent in prior experimental studies involving monolayer graphene samples with similar doubly-encapsulated geometry  Zibrov2018even ; Zeng2019 ; Chen2019 ; Polshyn2018corbino . It should be noted that LL mixing plays a significant role in facilitating interaction between composite fermions, creating favorable condition for pairing at half-filled $\Lambda$ levels  zhao2023composite ; wang2023next . Indeed, these unconventional states in GaAs were observed in hole-doped LL filling rather than electron-doped, presumably because the larger effective mass of holes, and consequently larger LL mixing, enhances interactions between CFs wang2022even ; wang2023next ; wang2023fractional . In quantum Hall graphene bilayers, we expect that the unique Coulomb environment reduces short-range repulsion in the CF sea, making it more favorable for CFs to form Cooper pairs. At large magnetic field, the absence of interlayer correlations restricts Cooper pairing to occur within each individual layer. As a result, the bilayer system offers a favorable platform to stabilize novel composite fermion states driven by intra-layer correlations, which are distinct from those observed in monolayer graphene samples.
In summary, we report a new type of bilayer exciton in quantum Hall graphene bilayers, constructed through pairing between quasiparticles and quasiholes of single-component FQHE states. At the Laughlin state, fractionally charged quasiparticles exhibit a braiding phase of $32\pi/$ 3. Because the particle and hole components of the exciton reside in separate layers, the resulting exciton inherits a total braiding phase of $34\pi/$ 3. This leads to a highly unusual composite particle that is overall charge-neutral but obeys anyonic quantum statistics. The flow dynamics of neutral anyons can be directly explored using counterflow transport, while their population can be uniquely controlled through layer-asymmetric field-effect doping. Together, this new type of bilayer exciton open a new frontier for engineering exotic quantum phases, poised to inspire broad future research directions.
acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. N.J.Z., R.Q.N., and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. N.J.Z. acknowledges partial support from the Jun-Qi fellowship. R.Q.N. and J.I.A.L. acknowledges partial support from the National Science Foundation EPSCoR Program under NSF Award OIA-2327206. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect those of the National Science Foundation. This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Alfred P. Sloan Foundation (G-2024-22395). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-1157490 and the State of Florida. N.B. and D.E.F. were supported in part by the National Science Foundation under Grant No. DMR-2204635. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.
Data availability
The data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
Author contribution
N.J.Z. and J.I.A.L. conceived the project. N.J.Z., R.Q.N., and X.L. fabricated the device. N.J.Z. and R.Q.N. performed the measurement. N.B. and D.E.F. provided theoretical inputs. K.W. and T.T. provided the material. N.J.Z., R.Q.N., N.B., D.E.F., and J.I.A.L wrote the manuscript together.
Competing financial interests
The authors declare no competing financial interests.
Supplementary Text
Supplementary Data
II Supplementary Text
II.1 Fractional charge and statistics of excitons belonging to ${}^{2}_{0}$ family
In this section we find the fractional charge and self-statistical phase for the Abelian interlayer excitons considered in the main text. We limit to the case when the effective filling $\nu_{i}^{*}\in\mathbb{Z}$ for each layer $i=1,2$. Since there are no interlayer fluxes, the bilayer ${\bf K}$-matrix is block diagonal ${\bf K}=\text{diag}({\bf K}_{1},{\bf K}_{2})$ with each block ${\bf K}_{i}$ being the monolayer ${\bf K}$-matrix along with their corresponding charge vector ${\bf t}_{i}$. Working in the symmetric representation ${\bf t}=({\bf t}_{1}\oplus{\bf t}_{2})^{T}=(1,\dots,1)$, each monolayer ${\bf K}$-matrix takes a simple form,
${\bf K}_{i}=\sigma_{i}{\bf I}+2{\bf J}~{}~{}~{}~{}\Leftrightarrow~{}~{}~{}~{}{\bf K}^{-1}_{i}=\sigma_{i}{\bf I}-\frac{2\sigma_{i}}{2\dim{\bf K}_{i}+\sigma_{i}}{\bf J}~{}~{}~{}~{}~{}~{}~{}\text{for}~{}~{}i=1,2,$ (S1)
where $\sigma_{i}=\pm 1$. The matrices ${\bf I}$ and ${\bf J}$ are identity and a constant matrix of ones respectively. Since $\dim{\bf K}_{i}$ represent the number of $U(1)$ gauge fields in the low energy Chern-Simons description, we can set $\dim{\bf K}_{i}=|\nu_{i}^{*}|$. Note that this identification only works for the Abelian states when $\nu_{i}^{*}\in\mathbb{Z}$ which we focus on in this section. Similarly, $\sigma_{i}$ is chosen depending upon the ${}^{2}_{0}$ particle-/hole-doped $\Lambda$ levels so we identify $\sigma_{i}=\text{sign}(\nu_{i}^{*})$.
Next, we compute the fractional charge for an arbitrary excitation labeled by ${\bf l}_{i}$ in the $i$th layer,
$Q_{i}=-e~{}{\bf t}^{T}_{i}{\bf K}_{i}^{-1}{\bf l}_{i}$ $=-e\sum_{n,m}^{|\nu_{i}^{*}|}\left(\sigma_{i}\delta_{nm}-\frac{2\sigma_{i}}{2|\nu_{i}^{*}|+\sigma_{i}}J_{nm}\right)\left[{\bf l}_{i}\right]_{m}=-e\left(\sigma_{i}-\frac{2\sigma_{i}}{2|\nu_{i}^{*}|+\sigma_{i}}|\nu_{i}|^{*}\right)\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$ (S2)
$=-e\frac{\ell_{i}}{\sigma_{i}\left(2\nu_{i}^{*}+1\right)},$ (S3)
where in the last equality we used $\sigma_{i}=\text{sign}(\nu_{i}^{*})$ and defined the integer $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$. Since the vector ${\bf l}_{i}$ has integer components, the minimal charge excitation in each layer is $Q_{\text{min.},i}=\pm e/|2\nu_{i}^{*}+1|$ which is simply the fundamental charge in the Jain states. Next, we find the minimal exciton charge for the bilayer system via the condition $Q_{1}+Q_{2}=0$ such that $Q_{1,2}\neq 0$. The minimal integer solutions of $\ell_{1},\ell_{2}\in\mathbb{Z}$ are,
$\frac{\ell_{1}}{|2\nu_{1}^{*}+1|}+\frac{\ell_{2}}{|2\nu_{2}^{*}+1|}=0~{}~{}~{}~{}\Rightarrow~{}~{}~{}~{}\ell_{i}=\frac{(-1)^{i}~{}|2\nu_{i}^{*}+1|}{\gcd(|2\nu_{i}^{*}+1|,|2\nu_{2}^{*}+1|)}~{}~{}~{}i=1,2,$ (S4)
where the integer valued function $\text{gcd}(n_{1},n_{2})$ computes the greatest common divisor of the integers $n_{1}$ and $n_{2}$. We also use $\text{sign}(\nu_{i}^{*})=\text{sign}(2\nu_{i}^{*}+1)$ since $|\nu_{i}^{*}|\geq 1$. This gives the minimal exciton charge,
$Q_{\text{exciton}}=\pm\frac{e}{\gcd(|2\nu_{i}^{*}+1|,|2\nu_{2}^{*}+1|)}.$ (S5)
The minimal solutions $\ell_{i}$ can also be used to compute the exciton statistics. Before we compute the exciton statistics, we notice that due to the definition $\ell_{i}\equiv\sum_{m}^{|\nu_{i}^{*}|}\left[{\bf l}_{i}\right]_{m}$, there is an ambiguity in defining the excitation corresponding to the integer vector ${\bf l}_{i}$. This ambiguity suggests that there are distinct solutions of ${\bf l}$ that all correspond to the same exciton charge, or for that matter, a generic quasiparticle charge. However, to make sure all these distinct solutions describe the same particle, we need to check if all such excitations result in the same mutual and self statistical phases modulo $2\pi$. We proved that the distinct solutions indeed result in the same statistical phases in our previous report (see Ref. Zhang2025fractionalexciton ) for a general bilayer ${}^{a}_{b}$ family. This guarantees that for any choice of ${\bf l}_{i}$ corresponding to the solutions obtained $\ell_{i}$ indeed describes the same minimal exciton as with any other choice of ${\bf l}_{i}$ preserving $\ell_{i}$. We thus set $\left[{\bf l}_{1}\right]_{1}=\ell_{1}$ and $\left[{\bf l}_{2}\right]_{1}=\ell_{2}$, and all the remaining elements are set to zero $[{\bf l}_{i}]_{n}=0$ $\forall~{}n\in[2,|\nu_{i}^{*}|]$. The exciton self-statistics is computed as,
$\theta_{s}$ $=\pi\sum_{i=1}^{2}{\bf l}_{i}^{T}{\bf K}_{i}^{-1}{\bf l}_{i}\mod 2\pi=\pi\left[{\bf K}_{1}^{-1}\right]_{11}\ell_{1}^{2}+\pi\left[{\bf K}_{2}^{-1}\right]_{11}\ell_{2}^{2}\mod 2\pi$ (S6)
$=\frac{\pi}{\gcd(|2\nu_{i}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}}\sum_{i=1}^{2}\left[\sigma_{i}-\frac{2}{(2\nu_{i}^{*}+1)}\right](2\nu_{i}^{*}+1)^{2}\mod 2\pi$ (S7)
$=\pi\sum_{i=1}^{2}\frac{\left(|2\nu_{i}^{*}+1|-2\right)\left(2\nu_{i}^{*}+1\right)}{\text{gcd}(|2\nu_{1}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}}\mod 2\pi$ (S8)
which shows that the excitons, in general, follow anyonic statistics rather than the usually anticipated bosonic character. Interestingly, we notice that since $\nu_{i}^{*}\in\mathbb{Z}$, the sum $2x\equiv\sum_{i=1}^{2}\left(|2\nu_{i}^{*}+1|-2\right)\left(2\nu_{i}^{*}+1\right)\in 2\mathbb{Z}$ for some integer $x\in\mathbb{Z}$. Similarly, the denominator $12y+1\equiv\text{gcd}(|2\nu_{1}^{*}+1|,|2\nu_{2}^{*}+1|)^{2}\in 2\mathbb{Z}+$ 1 for some $y\in\mathbb{Z}$. Now since the following equation has no integer solutions,
$\frac{x}{2y+1}\mod 1=\frac{1}{2}~{}~{}~{}~{}\Rightarrow~{}~{}~{}~{}\text{no solution for $x,y\in\mathbb{Z}$},$ (S9)
the excitons in the above construction are never fermionic for any combination of monolayer Jain states $\nu_{i}^{*}\in\mathbb{Z}$.
Figure S1: Activated behavior of quasiparticle excitations. (a) Arrhenius plot of bulk conductance measured in the parallel flow $G_{\text{PF}}$ (black) and counterflow $G_{\text{counter}}$ (blue) configurations at $\nu_{1}=\nu_{2}=1/3$ and $B$ = 28 T. The activated behavior of $G_{\text{PF}}$ is directly linked to the charge gap, $\Delta_{\text{charge}}$, which represents the energy required to create thermally excited quasiparticles and quasiholes. Interestingly, the activated behavior of $G_{\text{counter}}$ reveals the energy required to generate a charge-neutral excitation, $\Delta_{\text{exciton}}$, which closely aligns with $\Delta_{\text{charge}}$. This provides direct evidence that both quasiparticles and the neutral mode are generated through the same thermal process.